# Robust Equilibria in Shared Resource Allocation via Strengthening Border’s Theorem

### v1

@@ 5.2 Inducing Optimal Interim Allocation Probabilities
In this section, we show how to set the allocation probabilities $p_{i}^{S}$ to achieve optimal equilibrium performance in our BRB Mechanism. Specifically, in Theorem 4.2 we claimed that we can set each $p_{i}^{S}$ so that all interim allocation probabilities are equal: $p_{i}=1-\prod_{j}(1-\alpha_{j})$ for all $i$. This result follows from a special case of Border’s theorem, a theorem about the feasibility of reduced-form auctions. We present a special case of that theorem here and defer its full statement from the classical view of monetary auctions to Appendix C.
@@
Given arbitrary numbers $p_{i}\in[0,1]$, there exist allocation probabilities $p_{i}^{S}$ such that the $p_{i}$ are interim allocation probabilities induced by the $p_{i}^{S}$’s if and only if
$\sum_{i\in[n]}p_{i}\alpha_{i}=1-\prod_{i\in[n]}(1-\alpha_{i})$ (2)
and for every $I\subseteq[n]$,
$\sum_{i\in I}p_{i}\alpha_{i}\leq 1-\prod_{i\in I}(1-\alpha_{i}).$ (3)
@@
[…]
Figure 1: Flow network that can be used to prove Theorem 5.2. We let $S^{\prime}$ be the random set of bidding agents where each agent $i$ lies in $S^{\prime}$ independently with probability $\alpha_{i}$. Then, there are three kind of edges: edges whose flow corresponds to the probability of observing a specific $S^{\prime}$ (left), edges whose flow corresponds to how we randomly allocate the item condition on observing a specific $S^{\prime}$ (middle), and edges whose flow represent the probability that a specific agent gets the item (right). There is a flow of value $\Pr(S^{\prime}\neq\emptyset)$ if and only if there exist allocation probabilities $p_{i}^{S}$ inducing the interim allocation probabilities $p_{i}$. In other words, the flows $p_{i}^{S}\Pr(S^{\prime}=S)$ in the middle transform the probabilities that agents in a certain set $S$ bid to an agent $i\in S$ being allocated. We obtain the conditions in Theorem 5.2 by analyzing every minimum-cut of this network.
Now we can use Theorem 5.2 to prove Theorem 4.2.
Proof of Theorem 4.2.
We prove that the choice of $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$ satisfies the conditions of Theorem 5.2. Observe that (2) holds with this choice of $p_{i}$. Also, observe that (3) holds for $I=\emptyset$. To show that (3) holds for $I\neq\emptyset$, define $a_{I}$ as
$a_{I}=\frac{1-\prod_{i\in I}(1-\alpha_{i})}{\sum_{i\in I}\alpha_{i}}.$
With this choice of $p_{i}$, (3) says
This can be seen to be nonpositive by the identities that $1-x\leq e^{-x}$, so $\prod_{i\in I\setminus\{k\}}(1-\alpha_{i})\leq\exp\left(-\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)$, and also $e^{-y}(1+y)\leq 1$ applied to $y=\sum_{i\in I\setminus\{k\}}\alpha_{i}$. ∎
@@ 6 Robustness
In Section 5, we prove that each player playing an $\alpha_{i}$-aggressive strategy is an approximate equilibrium. To complete the proof of Theorem 4.3, we show how to choose the allocation probabilities $p_{i}^{S}$ to guarantee robustness while maintaining the same interim allocations. To do this, we strengthen Border’s Theorem to also satisfy some additional properties that guarantee robustness. We then show that this strengthening is necessary, as many allocations obtained by the standard Border’s Theorem are not robust. Finally, we show our robustness result is tight in that no matter the choice of $(p_{i}^{S})$, it is not possible to guarantee each agent a $\lambda$-robust strategy for $\lambda$ much greater than $1/2$.
@@ 6.1 Achieving a $1/2$ Robustness Factor
We will show how to choose allocation probabilities $p_{i}^{S}$ to make an $\alpha_{i}$-aggressive strategy $(\frac{1}{2}+\frac{1}{2}\alpha_{i}^{2})$-robust while still inducing the interim allocation probabilities $p_{i}=1-\prod_{k=1}^{n}(1-\alpha_{k})$ as in Theorem 4.2, proving Theorem 4.3.
We now consider (collusive) strategies that agents $j\neq i$ can employ to minimize agent $i$’s utility. Say that agent $i$ is blocked when she bids but does not receive the item. We will focus on how many bids the other agents have to make each time they block agent $i$. Let us first consider the case where other agents never bid two at a time. Conditioned on only agent $j$ bidding, agent $i$ gets blocked with probability $\alpha_{i}p_{j}^{\{i,j\}}$, with $\alpha_{i}$ being the probability that agent $i$ bids and $p_{j}^{\{i,j\}}$ being the probability that agent $j$ wins when they both bid. We have to ensure that this probability is not too large. Specifically, assume that for some $\bar{p}$ it holds $p_{j}^{\{i,j\}}\leq\bar{p}$ for all $j\neq i$. Thus, when one other agent bids at a time, the other agents must spend $(\bar{p}\alpha_{i})^{-1}$ tokens each time they block agent $i$ in expectation. When 2 or more agents bid, we still have that agent $i$ bids only with probability $\alpha_{i}$, so the expected number of tokens spent for blocking agent $i$ is at least $2\alpha_{i}^{-1}$. As long as $\bar{p}\geq 1/2$, this is less effective from a bang-per-buck perspective.
Since the number of times agents $j\neq i$ can bid is at most $\sum_{j\neq i}\alpha_{j}=(1-\alpha_{i})T$, the expected number of times agent $i$ can get blocked is at most $\max(\bar{p},1/2)\alpha_{i}(1-\alpha_{i})T$. Combining this with the fact that agent $i$ bids $\alpha_{i}T$ times in expectation, we get that she will win the item at least
@@
times in expectation. Therefore, to ensure that agent $i$ does not get too low utility, we have to ensure that $\bar{p}$ is not too large.
The formal statement of the above argument is the lemma below. Its proof, which we defer to Appendix D, uses martingale concentration arguments to ensure that agent $i$ obtains at least $(1-(1-\alpha_{i})\bar{p})$ fraction of her ideal utility with high probability.
@@
$\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]\geq\left(1-\bar{p}(1-\alpha_{i})\right)-O\left(\sqrt{\frac{\log T}{T}}\right)$
@@
We want to give allocation probabilities $p_{i}^{S}$ that induce interim allocation probabilities $p_{i}=1-\prod_{k}(1-\alpha_{k})$ that satisfy the above lemma with a low value of $\bar{p}$ to obtain simultaneous equilibrium and robustness guarantees as in Theorem 4.3. Let us now demonstrate that this is possible if each agent has equal fair share $\alpha_{i}=1/n$. Choose the symmetric probabilities $p_{i}^{S}=1/|S|$ for each $i\in S$, i.e., give the item to a bidding agent uniformly at random. By symmetry, these do indeed induce the interim allocation probabilities $p_{i}=1-\prod_{k}(1-\alpha_{k})=1-(1-1/n)^{n}$. They satisfy $p_{j}^{\{i,j\}}=1/2$, so we can use Lemma 6.1 with $\bar{p}=1/2$, yielding a robustness factor of $1-\bar{p}(1-\alpha_{i})=1/2+1/(2n)$.
It would be natural to try to generalize the use of $\bar{p}=1/2$ to the assymetric case. However, this would be too low, and the desired interim allocation of Theorem 4.3 may not be feasible with this restriction. To see why, it is useful to recall the two agent special case, when the agent with fair share $1-\alpha$ had to be allocated with probability $\alpha$, which can be more than $\bar{p}=1/2$. We’ll end up using this lemma with $\bar{p}=\frac{1+\alpha_{i}}{2}$ (always satisfying $\alpha_{i}\leq\bar{p}$), which proves the robustness guarantee of Theorem 4.3. The remaining technically challenging part of the proof is to show that there is a set of allocation probabilities $p_{i}^{S}$ satisfying these conditions.
The outline of the rest of the section is as follows. To prove that the desired allocation probabilities exist, consider the network of Fig. 1 used to prove Border’s theorem. Border’s theorem does not cover the upper bound constraint on $p_{j}^{\{i,j\}}$ of Lemma 6.1. To ensure this limit, we will change the capacity of the edges from node $u_{\{i,j\}}$ to node $v_{j}$ for all agents $j$ and all two-element sets $\{i,j\}$ to $\frac{1+\alpha_{i}}{2}$ as illustrated in Fig. 2.
In Lemma 6.2 we establish the condition for a flow of value $1-\prod_{k}(1-\alpha_{k})$ to exist with additional capacities on the edges that have infinite capacity in the original construction. While the condition is more complex, its proof is analogous to the proof of Border’s theorem discussed above: we need to make sure that each cut in the network has capacity high enough to support the desired flow amount. The most technically challenging and surprising part of our proof is establishing that the established condition actually holds in Lemma 6.3. This directly implies Theorem 4.3 completing the proof of the existence of allocation probabilities that simultaneously have optimal Nash equilibrium and the desired robustness.
While the below condition uses our upper bounds $p_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$ specifically, we can prove a more general condition with arbitrary upper bounds on any $p_{i}^{S}$, $S\subseteq[n],i\in S$. Because that condition is more complex, we present and prove it in Appendix C.
@@
There exists $p_{i}^{S}$’s that induce interim allocation probabilities $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$ and satisfy the upper bounds $p_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$ if and only if
$\left(1-\prod_{k=1}^{n}(1-\alpha_{k})\right)\sum_{i\in I}\alpha_{i}+\prod_{i\in I}(1-\alpha_{i})+\frac{1}{2}\prod_{k=1}^{n}(1-\alpha_{k})\sum_{i\in I}\frac{\alpha_{i}}{1-\alpha_{i}}\sum_{j\notin I}\alpha_{j}\leq 1$ (4)
for every $I\subseteq[n]$.
Proof.
@@
The flow network is depicted in Fig. 2.
[…]
Figure 2: Flow network used in the proof of Lemma 6.2. The flow network is very similar to the one in Fig. 1 used to prove Theorem 5.2. However, the red edges $(u_{\{i,j\}},v_{j})$ have explicit capacities as opposed to infinite capacity to enforce the additional bounds $\bar{p}_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$.
@@
$\sum_{i\in[n]}p_{i}\alpha_{i}=1-\prod_{i\in[n]}(1-\alpha_{i}).$ (5)
@@
using (5) for the last equality. Observe that in this flow network, allocation probabilities $(p_{i}^{S})$ satisfying upper bounds $p_{i}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$ induce the interim allocation probabilities $(p_{i})$ if and only if the flow $f$ is feasible where $f(s,u_{S})=c(s,u_{S})$, $f(v_{i},t)=c(v_{i},t)$, and $f(u_{S},v_{i})=p_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$. Since both the $s$-$t$ cuts with $s$ on one side and everything else on the other and the cut with $t$ on one side and everything else on the other both have cut capacity $1-\prod_{i\in[n]}(1-\alpha_{i})$, it suffices to show that (4) holds only if there is a feasible flow of flow value equal to this cut capacity.
@@
using (5) for the second-to-last equality and substituting $p_{i}=1-\prod_{k=1}^{n}(1-\alpha_{k})$ and doing some algebraic rearrangement for the last equality. Rearranging, (4) is equivalent to the above being least $1-\prod_{i\in[n]}(1-\alpha_{i})$. The result follows from the max-flow min-cut theorem. ∎
Next, we prove that allocation probabilities satisfying Lemma 6.2 do exist. The key observation to showing that (4) holds is noticing that the left-hand side of the inequality is Schur-concave in the variables $\alpha_{i}$ for $i\in I$ and Schur-convex in the variables $\alpha_{j}$ for $j\not\in I$. Using properties of Schur-convex and Schur-concave functions, it suffices to check the inequality for the special case when $\alpha_{i}$ is the same for all $i\in I$ and there is only one non-zero $\alpha_{j}$ outside of the set $I$. Fig. 3 plots the maximum possible value of the left hand side of (4), depending on the $\alpha$ values. We emphasize how close it gets to the bound of $1$ with sets $I$ of size $5$ or higher.
@@
$\left(1-\prod_{k=1}^{n}(1-\alpha_{k})\right)\sum_{i\in I}\alpha_{i}+\prod_{i\in I}(1-\alpha_{i})+\frac{1}{2}\prod_{k=1}^{n}(1-\alpha_{k})\sum_{i\in I}\frac{\alpha_{i}}{1-\alpha_{i}}\sum_{j\notin I}\alpha_{j}\leq 1.$ (6)
Proof.
@@
The left-hand side of (6) is precisely $f((\alpha_{i})_{i\in I},(\alpha_{j})_{j\notin I})$, so it suffices to show that the maximum of $f$ is at most $1$. By taking derivatives, we check in Appendix D that $f(\cdot,(y_{j})_{j\notin I})$ is Schur-concave for each $(y_{j})_{j\notin I}$ and that $f((x_{i})_{i\in I},\cdot)$ is Schur-convex for each $(x_{i})_{i\in I}$. Therefore, the maximum of $f$ occurs at $((x_{i}^{\star})_{i\in I},(y_{j}^{\star})_{j\notin I})$ when each $x_{i}^{\star}=X/m$ where $m=|S|$ and there is a single $y_{j^{\star}}^{\star}=1-X$ and all other $y_{j}^{\star}=0$. In that case,
@@
We plot $f$ in Fig. 3. It is easy to verify that $g(0)=1$, and hence if we set $X=0$ or $X=1$, then the above bound is $1$. To complete the proof, we claim that $g^{\prime}(X)\leq 0$ (i.e., $g(X)$ is decreasing) in $[0,1]$. This is indeed the case, as one can readily check
@@
Figure 3: Plot of $f(X)$, the maximum value of the left-hand side of (6) for a fixed size $m$ of $I$. It gets super close to $1$, and we prove that it never exceeds $1$.
Lemmas 6.2 and 6.3 establish the existence of allocation probabilities $p_{i}^{S}$ inducing interim allocation probabilities $p_{i}=1-\prod_{k=1}^{n}(1-\alpha_{k})$ that also satisfy upper bounds $p_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$. The interim allocation probabilities show the utility guarantee at equilibrium of Theorem 4.3. By Lemma 6.1 with $\bar{p}=\frac{1+\alpha_{i}}{2}$, the upper bounds prove the robustness claim of Theorem 4.3.

### v2

@@ 4.1 Inducing Optimal Interim Allocation Probabilities
Next we show how to set the allocation probabilities $p_{i}^{S}$ to achieve optimal equilibrium performance in our BRB Mechanism. Specifically, in Theorem˜3.5 we claimed that we can set each $p_{i}^{S}$ so that all interim allocation probabilities are equal: $p_{i}=1-\prod_{j}(1-\alpha_{j})$ for all $i$. This result follows from a special case of Border’s theorem on the feasibility of reduced-form auctions; we present this special case here and defer the full statement of the theorem to section˜D.
@@
Given arbitrary numbers $p_{i}\in[0,1]$, there exist allocation probabilities $p_{i}^{S}$ such that the $p_{i}$ are interim allocation probabilities induced by the $p_{i}^{S}$’s if and only if for every $I\subseteq[n]$,
(4.1) $\sum_{i\in I}p_{i}\alpha_{i}\leq 1-\prod_{i\in I}(1-\alpha_{i})$
with equality for $I=[n]$.
For completeness, we include a flow-based proof of Border’s theorem from [che2013generalized] in section˜D. Our special case (Theorem˜4.2) reduces the feasibility of interim allocations to a flow problem, illustrated in Figure˜4.1, where flows represent the probabilities of agent groups bidding and being allocated (see the figure caption for details). A feasible flow of value $\Pr(S^{\prime}\neq\emptyset)=1-\prod_{i\in[n]}(1-\alpha_{i})$ exists if and only if there are allocation probabilities $p_{i}^{S}$ that induce the given $p_{i}$. The necessity of condition (4.1) follows from interpreting the left-hand side as the total interim allocation to agents in set $I$, and the right-hand side as the probability that some agent in $I$ bids. Sufficiency follows by showing that these inequalities ensure every $(s,t)$ cut with finite capacity has at least the required flow value.
@@
Figure 4.1: Flow network that can be used to prove Theorem˜4.2. We let $S^{\prime}$ be the random set of bidding agents where each agent $i$ lies in $S^{\prime}$ independently with probability $\alpha_{i}$. Then, there are three kind of edges: edges whose flow corresponds to the probability of observing a specific $S^{\prime}$ (left), edges whose flow corresponds to how we randomly allocate the item condition on observing a specific $S^{\prime}$ (middle), and edges whose flow represent the probability that a specific agent gets the item (right). There is a flow of value $\Pr(S^{\prime}\neq\emptyset)$ if and only if there exist allocation probabilities $p_{i}^{S}$ inducing the interim allocation probabilities $p_{i}$. In other words, the flows $p_{i}^{S}\Pr(S^{\prime}=S)$ in the middle transform the probabilities that agents in a certain set $S$ bid to an agent $i\in S$ being allocated. We obtain the conditions in Theorem˜4.2 by analyzing every minimum-cut of this network.
In subsection˜5.1, we prove a generalization of Theorem˜4.2 and use it to prove a stronger result than Theorem˜3.5. That proof is quite involved, so for simplicity, we prove Theorem˜3.5 directly from Theorem˜4.2 in subsection˜E.2.
@@ 5 Robustness
In section˜4, we prove that each player playing an $\alpha_{i}$-aggressive strategy is an approximate equilibrium. To complete the proof of Theorem˜3.6, we show how to choose the allocation probabilities $p_{i}^{S}$ to guarantee robustness while maintaining the same interim allocations. To do this, in subsection˜5.1, we strengthen Border’s Theorem to also satisfy some additional properties that guarantee $1/2$-robustness. In subsection˜5.2, we then show that this strengthening is necessary, as many allocations obtained by the standard Border’s Theorem are not robust. Finally, in subsection˜5.3, we show our robustness result is tight in that no matter the choice of $(p_{i}^{S})$, in the BRB mechanism, it is not possible to guarantee each agent a $\lambda$-robust strategy for a constant $\lambda$ greater than $1/2$ not depending on the fair shares $(\alpha_{i})$ or the time horizon $T$.
@@ 5.1 Achieving a $1/2$ Robustness Factor
To prove Theorem˜3.6, we show how to choose probabilities $p_{i}^{S}$ to make the $\alpha_{i}$-aggressive strategy $(\frac{1}{2}+\frac{1}{2}\alpha_{i}^{2})$-robust while still inducing interim allocation probabilities $p_{i}=1-\prod_{k=1}^{n}(1-\alpha_{k})$ (as in Theorem˜3.5).
We consider (collusive) strategies that agents $j\neq i$ can employ to minimize agent $i$’s utility. Say that agent $i$ is blocked when she bids but does not receive the item. We will focus on how many bids the other agents have to make each time they block agent $i$. Let us first consider the case where other agents never bid two at a time. Conditioned on only agent $j$ bidding, agent $i$ gets blocked with probability $\alpha_{i}p_{j}^{\{i,j\}}$, with $\alpha_{i}$ being the probability that agent $i$ bids and $p_{j}^{\{i,j\}}$ being the probability that agent $j$ wins when they both bid. We have to ensure that this probability is not too large. Specifically, assume that for some $\bar{p}$ it holds $p_{j}^{\{i,j\}}\leq\bar{p}$ for all $j\neq i$. Thus, when one other agent bids at a time, the other agents must spend $(\bar{p}\alpha_{i})^{-1}$ tokens each time they block agent $i$ in expectation. When 2 or more agents bid, we still have that agent $i$ bids only with probability $\alpha_{i}$, so the expected number of tokens spent for blocking agent $i$ is at least $2\alpha_{i}^{-1}$. As long as $\bar{p}\geq 1/2$, this is less effective from a bang-per-buck perspective.
Since the number of times agents $j\neq i$ can bid is at most $\sum_{j\neq i}\alpha_{j}=(1-\alpha_{i})T$, the expected number of times agent $i$ can get blocked is at most $\max(\bar{p},1/2)\alpha_{i}(1-\alpha_{i})T$. Combining this with the fact that agent $i$ bids $\alpha_{i}T$ times in expectation, we get that the expected number of rounds in which she wins the item is at least
@@
To ensure that agent $i$ does not get too low utility, we thus need that $\bar{p}$ is not too large.
The formal statement of the above argument is the lemma below. Its proof, which we defer to subsection˜E.5, uses martingale concentration arguments to ensure that agent $i$ obtains at least $(1-(1-\alpha_{i})\bar{p})$ fraction of her ideal utility with high probability.
@@
$\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]\geq\left(1-\bar{p}(1-\alpha_{i})\right)v_{i}^{\star}-O\left(\sqrt{\frac{\log T}{T}}\right)$
@@
We want to give allocation probabilities $p_{i}^{S}$ that induce interim probabilities $p_{i}=1-\prod_{k}(1-\alpha_{k})$, while keeping the upper bounds $\bar{p}$ as in Lemma˜5.1 small. This enables us to achieve both equilibrium and robustness guarantees, as stated in Theorem˜3.6.
When all agents have equal fair share $\alpha_{i}=1/n$, assigning the item uniformly at random among bidders (i.e., $p_{i}^{S}=1/|S|$) yields the desired interim allocation $p_{i}=1-(1-1/n)^{n}$ by symmetry, and satisfies $p_{j}^{\{i,j\}}=1/2$. Applying Lemma˜5.1 with $\bar{p}=1/2$ then gives a robustness guarantee of $1/2+1/(2n)$.
It would be natural to try to generalize the use of $\bar{p}=1/2$ to the asymmetric case. However, this would be too low, and the desired interim allocation of Theorem˜3.6 may not be feasible with this restriction. In fact, in section˜A, we show that if there are only $2$ agents, to induce the desired interim allocation probabilities, the unique solution is $p_{1}^{\{1,2\}}=\alpha_{2}$, which can be more than $\bar{p}=1/2$. We end up using $\bar{p}=\frac{1+\alpha_{i}}{2}$ (satisfying $\alpha_{i}\leq\bar{p}$), which proves the robustness guarantee of Theorem˜3.6. The remaining technically challenging part of the proof is to show that there is a set of allocation probabilities $p_{i}^{S}$ satisfying these conditions.
To prove that the desired allocation probabilities exist, we consider the network of Figure˜4.1 used to prove Border’s theorem. Border’s theorem does not cover upper bound constraints on $p_{j}^{\{i,j\}}$ of Lemma˜5.1. To ensure this limit, we change the capacity of the edges from node $u_{\{i,j\}}$ to node $v_{j}$ for all agents $j$ and all two-element sets $\{i,j\}$ to $\frac{1+\alpha_{i}}{2}$ as illustrated in Figure˜5.1. In Lemma˜5.2, we establish conditions under which a flow of value $1-\prod_{k}(1-\alpha_{k})$ exists in this modified network. While the condition is more complex, the proof mirrors the original Border argument: we verify that every cut has sufficient capacity to support the desired flow. The most technically challenging and surprising part is establishing that this condition actually holds, which we do in Lemma˜5.3, thereby completing the proof of Theorem˜3.6.
Finally, while our condition uses specific bounds $p_{j}^{\{i,j\}}\leq(1+\alpha_{i})/2$, we also prove a more general version allowing arbitrary upper bounds on any $p_{i}^{S}$. This generalization is presented in section˜D.
[…]
Figure 5.1: Flow network for the proof of Lemma˜5.2, similar to the one in Figure˜4.1 used in Theorem˜4.2. However, the red edges $(u_{\{i,j\}},v_{j})$ have explicit capacities as opposed to infinite capacity to enforce the additional bounds $\bar{p}_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$.
@@
There exists $p_{i}^{S}$’s that induce interim allocation probabilities $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$ and satisfy the upper bounds $p_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$ if and only if for every $I\subseteq[n]$,
(5.1) $\left(1-\prod_{k=1}^{n}(1-\alpha_{k})\right)\sum_{i\in I}\alpha_{i}+\prod_{i\in I}(1-\alpha_{i})+\frac{1}{2}\prod_{k=1}^{n}(1-\alpha_{k})\sum_{i\in I}\frac{\alpha_{i}}{1-\alpha_{i}}\sum_{j\notin I}\alpha_{j}\leq 1.$
Proof 1.
@@
The flow network is depicted in Figure˜5.1.
@@
(5.2) $\sum_{i\in[n]}p_{i}\alpha_{i}=1-\prod_{i\in[n]}(1-\alpha_{i}).$
@@
using (5.2) for the last equality. Observe that in this flow network, allocation probabilities $(p_{i}^{S})$ satisfying upper bounds $p_{i}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$ induce the interim allocation probabilities $(p_{i})$ if and only if the flow $f$ is feasible where $f(s,u_{S})=c(s,u_{S})$, $f(v_{i},t)=c(v_{i},t)$, and $f(u_{S},v_{i})=p_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$. Since both the $s$-$t$ cuts with $s$ on one side and everything else on the other and the cut with $t$ on one side and everything else on the other both have cut capacity $1-\prod_{i\in[n]}(1-\alpha_{i})$, it suffices to show that (5.1) holds only if there is a feasible flow of flow value equal to this cut capacity.
@@
using (5.2) for the second-to-last equality and substituting $p_{i}=1-\prod_{k=1}^{n}(1-\alpha_{k})$ and doing some algebraic rearrangement for the last equality. Rearranging, (5.1) is equivalent to the above being least $1-\prod_{i\in[n]}(1-\alpha_{i})$. The result follows from the max-flow min-cut theorem.
Next, we prove that allocation probabilities satisfying Lemma˜5.2 do exist. The key observation to showing that (5.1) holds is noticing that the left-hand side of the inequality is Schur-concave in the variables $\alpha_{i}$ for $i\in I$ and Schur-convex in the variables $\alpha_{j}$ for $j\not\in I$. Using properties of Schur-convex and Schur-concave functions, it suffices to check the inequality for the special case when $\alpha_{i}$ is the same for all $i\in I$ and there is only one non-zero $\alpha_{j}$ outside of the set $I$. Figure˜5.2 plots the maximum possible value of the left hand side of (5.1), depending on the $\alpha$ values. We emphasize how close it gets to the bound of $1$ with sets $I$ of size $5$ or higher.
@@
(5.3) $\left(1-\prod_{k=1}^{n}(1-\alpha_{k})\right)\sum_{i\in I}\alpha_{i}+\prod_{i\in I}(1-\alpha_{i})+\frac{1}{2}\prod_{k=1}^{n}(1-\alpha_{k})\sum_{i\in I}\frac{\alpha_{i}}{1-\alpha_{i}}\sum_{j\notin I}\alpha_{j}\leq 1.$
Proof 2.
@@
The left-hand side of (5.3) is precisely $f((\alpha_{i})_{i\in I},(\alpha_{j})_{j\notin I})$, so it suffices to show that the maximum of $f$ is at most $1$. By taking derivatives, we check in subsection˜E.6 that $f(\cdot,(y_{j})_{j\notin I})$ is Schur-concave for each $(y_{j})_{j\notin I}$ and that $f((x_{i})_{i\in I},\cdot)$ is Schur-convex for each $(x_{i})_{i\in I}$. Therefore, the maximum of $f$ occurs at $((x_{i}^{\star})_{i\in I},(y_{j}^{\star})_{j\notin I})$ when each $x_{i}^{\star}=X/m$ where $m=|S|$ and there is a single $y_{j^{\star}}^{\star}=1-X$ and all other $y_{j}^{\star}=0$. In that case,
@@
We plot $f$ in Figure˜5.2. It is easy to verify that $g(0)=1$, and hence if we set $X=0$ or $X=1$, then the above bound is $1$. To complete the proof, we claim that $g^{\prime}(X)\leq 0$ (i.e., $g(X)$ is decreasing) in $[0,1]$. This is indeed the case, as one can readily check
@@
Figure 5.2: Plot of $f(X)$, the maximum value of the left-hand side of (5.3) for a fixed size $m$ of $I$. It gets super close to $1$, and we prove that it never exceeds 11.
Lemmas˜5.2 and 5.3 establish the existence of allocation probabilities $p_{i}^{S}$ inducing interim allocation probabilities $p_{i}=1-\prod_{k=1}^{n}(1-\alpha_{k})$ that also satisfy upper bounds $p_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$. The interim allocation probabilities show the utility guarantee at equilibrium of Theorem˜3.6. By Lemma˜5.1 with $\bar{p}=\frac{1+\alpha_{i}}{2}$, the upper bounds prove the robustness claim of Theorem˜3.6.